# Black Hole Entropy, Quantum Corrections and EFT Transitions

### v1

@@ Acknowledgements
We are indebted to José Calderón-Infante, Álvaro Herráez, Luis Ibáñez, Dieter Lüst, Sameer Murthy, Tomás Ortín, Eran Palti, Alejandro Ruipérez, Savdeep Sethi and Ángel Uranga for illuminating discussions and very useful comments on the manuscript. We also acknowledge valuable conversations with Alek Bedroya, Jake McNamara, Cumrun Vafa and Max Wiesner. A.C. would like to thank José Calderón-Infante and Álvaro Herráez for collaboration on related topics. The authors thank IFT-Madrid for hospitality and support during the different stages of this work. A.C. also acknowledges the hospitality of the Department of Physics at Harvard University during the early stages of this work. The work of A.C. is supported by a Kadanoff as well as an Associate KICP fellowships. A.C. and M.Z. are also grateful to Teresa Lobo and Miriam Gori for their continuous encouragement and support.
@@ A.1 Asymptotic expansions and optimal truncation
Notice that the above definition resembles —but is actually different than— that corresponding to convergent power series (i.e., Taylor/Laurent expansions). Hence, the infinite series specified by (A.1) might be non-convergent, but nonetheless it must be that condition (A.2) holds for any of its finite order truncations. In fact, this concept can be readily extended to include complex functions as well, even though in that case one usually needs to be slightly more careful about the validity regime of the approximation due to e.g., Stokes’ phenomena Dorigoni:2014hea (see also the discussion around eq. (A.14) below). Furthermore, it is easy to show that a function can have at most one asymptotic expansion around some point $x_{0}$ (or infinity, see footnote 29), but the reverse statement is not true, i.e., two different functions $f(x)$ and $h(x)$ may share the same asymptotic series at a given point within their domain of definition.
@@ B.1 The Taub-NUT geometry
Figure 10: Schematic depiction of the Taub-NUT geometry for a single Kaluza-Klein (KK) monopole. It can be regarded as a gravitational instanton interpolating between an asymptotic spacetime with local topology $\mathbb{R}^{3}\times\mathbf{S}^{1}$ and a smooth configuration at its core locally of the form $\mathbb{R}^{4}$. Asymptotically, the circle radius $R$ is much smaller than the radial coordinate $r\gg R$. Close to the core, the circle radius $R_{\epsilon}$ scales as the radial coordinate $R_{\epsilon}\sim\epsilon$ and the circle fibration of $\mathbf{S}^{1}$ on the $\mathbf{S}^{2}\subset\mathbb{R}^{3}$ has the topological structure of an $\mathbf{S}^{3}$. For $N$ KK monopoles the sphere is replaced by the lens space $\mathbf{S}^{3}/\mathbb{Z}_{N}$.
[…]
@@
$N$-coincident Kaluza-Klein monopoles
@@
[…]
@@ B.2.1 The 4d Perspective
which prompts us to identify the 5d and 4d central charges as follows (recall that $8\pi G_{4}=\kappa_{4}^{2}$) [^36] This can also be deduced directly from the map relating the 4d and 5d gauge charges, cf. footnote 21.
@@ B.2.2 Connecting the 4d and 5d pictures
[…]

### v2

@@ Acknowledgements
We are indebted to José Calderón-Infante, Matilda Delgado, Álvaro Herráez, Luis Ibáñez, Dieter Lüst, Sameer Murthy, Tomás Ortín, Eran Palti, Alejandro Ruipérez, Savdeep Sethi and Ángel Uranga for illuminating discussions and useful comments on the manuscript. We also acknowledge valuable conversations with Alek Bedroya, Jake McNamara, Cumrun Vafa and Max Wiesner. A.C. would like to thank José Calderón-Infante and Álvaro Herráez for collaboration on related topics. The authors thank IFT-Madrid for hospitality and support during the different stages of this work. A.C. acknowledges the hospitality of the Department of Physics at Harvard University during the early stages of this work. The work of A.C. is supported by a Kadanoff and an Associate KICP fellowships, as well as through the NSF grants PHY-2014195 and PHY-2412985. A.C. and M.Z. are also grateful to Teresa Lobo and Miriam Gori for their continuous encouragement and support.
@@ A.1 Asymptotic expansions and optimal truncation
Notice that the above definition resembles —but is actually different than— that corresponding to convergent power series (i.e., Taylor/Laurent expansions). Hence, the infinite series specified by (A.1) might be non-convergent, but nonetheless it must be that condition (A.2) holds for any of its finite order truncations. In fact, this concept can be readily extended to include complex functions as well, even though in that case one usually needs to be slightly more careful about the validity regime of the approximation due to e.g., Stokes’ phenomena Dorigoni:2014hea (see also the discussion around eq. (A.14) below). Furthermore, it is easy to show that a function can have at most one asymptotic expansion around some point $x_{0}$ (or infinity, see footnote 31), but the reverse statement is not true, i.e., two different functions $f(x)$ and $h(x)$ may share the same asymptotic series at a given point within their domain of definition.
@@ B.1 The Taub-NUT geometry
Figure 10: Schematic depiction of the Taub-NUT geometry for a single Kaluza-Klein (KK) monopole. It can be regarded as a gravitational instanton interpolating between an asymptotic spacetime with local topology $\mathbb{R}^{3}\times\mathbf{S}^{1}$ and a smooth configuration at its core locally of the form $\mathbb{R}^{4}$. Asymptotically, the circle radius $R$ is much smaller than the radial coordinate $r\gg R$. Close to the core, the circle radius $R_{\epsilon}$ scales as the radial coordinate $R_{\epsilon}\sim\epsilon$ and the circle fibration of $\mathbf{S}^{1}$ on the $\mathbf{S}^{2}\subset\mathbb{R}^{3}$ has the topological structure of an $\mathbf{S}^{3}$. For NN KK monopoles the sphere is replaced by the lens space $\mathbf{S}^{3}/\mathbb{Z}_{N}$.
are left-invariant [^35] The right-invariant set may be obtained directly from (B.2) upon interchanging $\psi\leftrightarrow\phi$. 1-forms of $SU(2)\cong\mathbf{S}^{3}$ defined in terms of the Euler angles, namely $0\leq\theta<\pi,\,0\leq\phi<2\pi,\,0\leq\psi<4\pi$. These covectors moreover satisfy the algebra relations
@@
NN-coincident Kaluza-Klein monopoles
@@
with $d\tilde{\Omega}_{3}^{2}=\frac{1}{4}(d\Omega_{2}^{2}+\left(d\psi/p^{0}+\cos\theta d\phi\right)^{2})$ denoting the metric on $\mathbf{S}^{3}/\mathbb{Z}_{p^{0}}$. Hence, for the particular case of black holes having $J_{\rm L}=0$, we deduce that the horizon radius behaves as $r_{h,\,5d}^{2}=(2/\pi)^{2/3}p^{0}|Z_{\rm 5d}|G_{5}^{2/3}$, thus providing an entropy of the form
$\mathcal{S}_{\rm BH}=\frac{A_{\rm hor}}{4G_{5}}=\frac{2\pi^{2}r_{h,\,5d}^{3}}{4p^{0}G_{5}}=\pi\sqrt{p^{0}|Z_{\rm 5d}|^{3}}\,$, (B.18)
where the additional $1/p^{0}$ factor comes from the reduced volume of $\mathbf{S}^{3}/\mathbb{Z}_{p^{0}}$ with respect to that of the round 3-sphere. Notice that, for large central charges, the BPS solution (B.13) should be more accurately regarded as a four-dimensional black hole, since its radius is larger than the asymptotic value of the $\mathbf{S}^{1}$ component. Conversely, if $r_{h,\,5d}$ is small compared to the circle at infinity, it must be seen as a 5d black hole at the center of a Taub-NUT geometry. Therefore, assuming we are in the second scenario, one may ask about the relative size of the black hole horizon and the compact $\mathbf{S}^{1}$, when evaluated at $r=0$. The latter can be easily determined from eqs. (B.13) and (B.15) to be $\rho_{5}^{2}=(2/\pi)^{2/3}\,|Z_{\rm 5d}|G_{5}^{2/3}/p^{0}$, which implies that
$\frac{\rho_{5}}{r_{h,\,5d}}=\frac{1}{p^{0}}\,.$ (B.19)
@@ B.2.1 The 4d Perspective
which prompts us to identify the 5d and 4d central charges as follows (recall that $8\pi G_{4}=\kappa_{4}^{2}$) [^38] This can also be deduced directly from the map relating the 4d and 5d gauge charges, cf. footnote 23.
@@ B.2.2 Connecting the 4d and 5d pictures
Before we conclude this appendix, we would like to emphasize several interesting features that emerge from the discussion above. First, one might wonder why we have been careful to distinguish between the five-dimensional ratio $\rho_{5}/r_{h,\,5d}$ (cf. (B.19)) and the analogous one (i.e., $r_{5}/r_{h,\,4d}$) computed in the 4d supergravity theory, see eq. (B.32). The key point here is that, since the 4d and 5d entropies match, the corresponding black hole radii must necessarily differ. More precisely, they satisfy
$\frac{4\pi r_{h,\,4d}^{2}}{4(G_{5}/4\pi R)}=\frac{2\pi^{2}r_{h,\,5d}^{3}}{4p^{0}G_{5}}\,$, (B.33)